A smooth sphere A of mass 3m moving on a smooth horizontal table with speed 4u, impiges directly on another smooth sphere B of mass 2m, moving with speed u in the opposite direction to A. The coefficient of restitution between A and B is e.
(a) Find in terms of e and u, the speed of B after the impact.
(b) Show that e = 1/3
(c) Find the impulse exerted on A by the impact.
1
Expert's answer
2021-02-07T17:23:40-0500

"\begin{cases}
 12mu-2mu=3mx+2my,\\
 3m\cdot 16u^2+2mu^2=3mx^2+2my^2;
\end{cases}" "\begin{cases}
 10u=3x+2y,\\
 50u^2=3x^2+2y^2;
\end{cases}"

"\begin{cases}
 100u^2=9x^2+12xy+4y^2,\\
 100u^2=6x^2+4y^2;
\end{cases}"

"0=3x^2+12xy,"

"3x(x+4y)=0,"

"x=-4y,"

"-10y=10u,"

"y=-u,"

"x=4u."